OpenStudy (anonymous):

value of logarithmic expressions: log2 (8)

OpenStudy (lgbasallote):

hint \[\huge \log _a b \implies \log b \div \log a\] does that help?

OpenStudy (anonymous):

4

OpenStudy (unklerhaukus):

\[\log_28=\log_22^3=\]

OpenStudy (lgbasallote):

really?

OpenStudy (lgbasallote):

how'd you get 4?